Solve the given equation.

4a^(2/3)-16=0
1
Expert's answer
2013-01-23T08:58:19-0500
4a^(2/3)-16=0
4a^(2/3)=16
a^(2/3)=4
(a^(1/3))^2=4
so a^(1/3)=-2 or a^(1/3)=2
a=-8 or a=8